OpenStudy (anonymous):

Find the second term in this sequence: t sub n=(6-n)^2

OpenStudy (anonymous):

Well, n is the term. n = (6-2)^2 n = (4)^2 n = 16

OpenStudy (anonymous):

This means the 2nd term will be 16.

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

No problem good luck (: